OpenStudy (adamaero):

How does the integral[-sin(x)*cos^2(x)]dx equal (1/3)cos^3(x) + C?

OpenStudy (adamaero):

\[-\int\limits_{}^{} \sin(x)*\cos^2(x)dx = \frac{ 1 }{ 3 } \cos^3(x) + C\] I got a longer result.

OpenStudy (freckles):

did you try doing letting u=cos(x)?

OpenStudy (freckles):

the answer can look different by the way and still represent the same family of functions

OpenStudy (freckles):

but anyways the letting of u=cos(x) will lead to that exact result you are asking about